Determine the hydrostatic force on a triangular plate that is submerged in water.

6m wide and 4m high.


1
Expert's answer
2021-05-12T07:51:48-0400


hydrostatic force = pressure "\times Area"


pressure = "\rho gdy"


= "1000 \times 9.8 \times dy"


from similar triangles


"\frac{6}{w}= \frac{4}{y}"


"w= \frac{3}{2}y"


force therefore = "1000\times 9.8\times (4-y)(\frac{3}{2})dy"

= "\int_0^4 9800\times (4-y)(\frac{3}{2})dy"


hydrostatic force = "156800 N"